# Set Features for Anomaly Detection

### v2

@@ Abstract
This paper proposes set features for detecting anomalies in samples that consist of unusual combinations of normal elements. Many leading methods discover anomalies by detecting an unusual part of a sample. For example, state-of-the-art segmentation-based approaches, first classify each element of the sample (e.g., image patch) as normal or anomalous and then classify the entire sample as anomalous if it contains anomalous elements. However, such approaches do not extend well to scenarios where the anomalies are expressed by an unusual combination of normal elements. In this paper, we overcome this limitation by proposing set features that model each sample by the distribution of its elements. We compute the anomaly score of each sample using a simple density estimation method, using fixed features. Our approach [^1] The code is available on github: https://github.com/NivC/SINBAD outperforms the previous state-of-the-art in image-level logical anomaly detection and sequence-level time series anomaly detection.
@@ 3.3 Set Features by Histogram of Projections
Figure 3: For both image and time series samples, we extract set elements at different granularity. For images (left), the sets of elements are extracted from different ResNet levels. For time series (right), we take pyramids of windows at different strides around each time step (noted in blue circles).
@@ 5.1 Logical Anomaly Detection Results
The anomalies in each class are divided into structural anomalies and logical anomalies. Structural anomalies feature local defects, somewhat similar to previous datasets such as [3]. Conversely, logical anomalies may violate ‘logical’ conditions expected from the normal data. As one example, an anomaly may include a different number of objects than the numbers expected from a normal sample (while all the featured object types exist in the normal class (Fig.1)). Other types of logical anomalies in the dataset may include cases where distant parts of an image must correlate with one another. For instance, within the normal data, the color of one object may correlate with the length of another object. These correlations may break in an anomalous sample.
Baselines. We compare to baseline methods used by the paper which presented the MVTec-LOCO dataset [5]: Variational Model (VM) [56], MNAD, f-AnoGAN [54], AE / VAE. Student Teacher (ST), SPADE, PatchCore (PCore) [47]. We also compare to GCAD [5] - a reconstruction-based method, based on both local and global deep ResNet features, which was explicitly designed for logical anomaly detection. A concurrent method, EfficientAD [1], focuses on structural anomalies and achieves impressive results there (but underperforms on logical anomalies). It is not included in our table as per-class results are not reported.
@@
Results. We report our results on image-level detection of logical anomalies and structural anomalies in Tab.1. Interestingly, we find complementary strengths between our approach and GCAD, a reconstruction-based approach by [5]. Although GCAD performed better on specific classes (e.g., pushpins), our approach provides better results on average. Notably, our approach provides non-trivial anomaly detection capabilities on the screw bag class, while baseline approaches are close to the random baseline. All other approaches performed significantly worse on the logical anomaly classes.
@@ 5.4 Ablations
Using individual ResNet levels. In Tab.4 we report the results of our method when different components of our multi-level ResNet ensemble are removed. We report the results using only the representation from the third or fourth ResNet block (“Only 3 / 4”). We also report the results of using both ResNet blocks but without the raw-pixels level (“No Pixels”).
@@
Ablating our histogram density-estimation method. In Tab.3 we ablate different aspects of our use of histogram set descriptors. Simple averaging. We compare to a simple averaging [35] of the set features (Fig. 2), ablating our entire set-features approach. This yields a significantly worse performance. No random projection. We ablate our use of random projections as described in Sec.3.3. We replace the random histograms with similar histograms using the raw given features. No whitening. We ablate our Gaussian modelling of the set features. The whitening is not essential for the image modality, as it is for the time-series data (Table 12).
@@
Ablating the number of bins and the number of projections. While generally we would like to have as many random projections as possible; and a large number of bins per histogram (as long we have enough statistics to estimate the occupancy in each of them) we find that in practice the values we choose are large enough. We show in the supplementary material (App. Tab.10,11) that while significantly lower values in these parameters degrade our performance, the benefit from using larger values saturates.
@@ 7 Limitations
Class-specific performance. While our method outperforms on average, in some classes we do not perform as well compared to baseline approaches. A better understanding of the cases when our method fails would be beneficial for deploying it in practice.
@@ 15 Further Image Anomaly Detection Ablation
We include here the ablation tables for the number of random projections and number of bins, for logical anomaly detection (Tab.10,11).
@@ 16 Time Series Anomaly Detection Ablations
Effect of number of pyramid levels and window size. We ablate the two hyperparameters of the time-series feature extraction: the number of pyramid windows used $L$, and the number of samples per window $\tau$ (see Sec.4.2). We find that in both cases the results are not sensitive to the chosen parameters (Tab.14,15).

### v3

@@ Abstract
This paper proposes to use set features for detecting anomalies in samples that consist of unusual combinations of normal elements. Many leading methods discover anomalies by detecting an unusual part of a sample. For example, state-of-the-art segmentation-based approaches, first classify each element of the sample (e.g., image patch) as normal or anomalous and then classify the entire sample as anomalous if it contains anomalous elements. However, such approaches do not extend well to scenarios where the anomalies are expressed by an unusual combination of normal elements. In this paper, we overcome this limitation by proposing set features that model each sample by the distribution of its elements. We compute the anomaly score of each sample using a simple density estimation method, using fixed features. Our approach [^1] The code is available on github: https://github.com/NivC/SINBAD outperforms the previous state-of-the-art in image-level logical anomaly detection and sequence-level time series anomaly detection.
@@ 3.3 Set Features by Histogram of Projections
Figure 3: For both image and time series samples, we extract set elements at different granularity. For images (left), the sets of elements are extracted from different ResNet levels. For time series (right), we take pyramids of windows at different strides around each time step.
@@ 5 Results
[…]
SINBAD EfficientAD (reported) EfficientAD (reproduced) PUAD SINBAD+EfficientAD
Logical 91.2 $\pm$ 0.8 86.8 85.9 $\pm$ 0.4 92.0 92.7 $\pm$ 0.6
Structural 85.5 $\pm$ 0.7 94.7 93.8 $\pm$ 0.4 94.1 95.8 $\pm$ 0.5
All 88.3 $\pm$ 0.7 90.8 89.8 $\pm$ 0.5 93.1 94.2 $\pm$ 0.6
@@ 5.1 Logical Anomaly Detection Results
The anomalies in each class are divided into structural anomalies and logical anomalies. Structural anomalies feature local defects, somewhat similar to previous datasets such as [3]. Conversely, logical anomalies may violate ‘logical’ conditions expected from the normal data. As one example, an anomaly may include a different number of objects than the numbers expected from a normal sample (while all the featured object types exist in the normal class Fig.1). Other types of logical anomalies in the dataset may include cases where distant parts of an image must correlate with one another. For instance, within the normal data, the color of one object may correlate with the length of another object. These correlations may break in an anomalous sample.
Baselines. We compare to baseline methods used by the paper which presented the MVTec-LOCO dataset [5]: Variational Model (VM) [57], MNAD, f-AnoGAN [55], AE / VAE. Student Teacher (ST), SPADE, PatchCore (PCore) [48]. We also compare to GCAD [5] - a reconstruction-based method, based on both local and global deep ResNet features, which was explicitly designed for logical anomaly detection. EfficientAD - [1], a reconstruction-based method, with a loss aimed at preventing an autoencoder from reconstructing well unseen images. We also report the results by PUAD [58], an ensemble method combining [1] and [47]. Finally, we report SINBAD+EfficientAD, a simple average of our and EfficientAD’s per-sample results. As the last set of baselines does not always report per-class accuracies, we report them in a different table.
@@
Results. We report per-class results on image-level detection of logical anomalies and structural anomalies in Tab.1. Interestingly, we find complementary strengths between our approach and GCAD, a reconstruction-based approach by [5]. Although GCAD performed better on specific classes (e.g., pushpins), our approach provides better results on average. Notably, our approach provides non-trivial anomaly detection capabilities on the screw bag class, while baseline approaches are close to the random baseline. EfficientAD [1], focuses on structural anomalies and achieves impressive results on them, but underperforms on logical anomalies (see Tab.2).
As our method is comparatively strong on specific classes (e.g., Screw bag Logical, $81.1\%$ compared to $56.0\%$ of GCAD and $55.5\%$ of EfficientAD), it makes a strong contributor to ensemble methods. For example, a simple combination of our method with EfficientAD [1] (SINBAD+EfficientAD) outperforms all other methods and ensembles (See Tab. 2).
@@ 5.4 Ablations
Using individual ResNet levels. In Tab.5 we report the results when different components of our multi-level ResNet ensemble are removed. We report the results using only the representation from the third or fourth ResNet block (“Only 3 / 4”). We report the results of using both ResNet blocks but without the raw-pixels level (“No Pixels”).
@@
Ablating our histogram density-estimation method. In Tab.4 we ablate different aspects of our histogram set descriptors. Simple averaging. We show a simple averaging [35] of the set features (Fig. 2), ablating our entire set-features approach. This yields a significantly worse performance. No random projection. We ablate our use of random projections (Sec.3.3). We replace the random histograms with similar histograms using the raw given features. No whitening. We ablate our Gaussian modelling of the set features. The whitening is not essential for the image modality, as it is for the time-series data (Table 13).
@@
Ablating the number of bins and the number of projections. While generally we would like to have as many random projections as possible; and a large number of bins per histogram (as long we have enough statistics to estimate the occupancy in each of them) we find that in practice the values we choose are large enough. We show in the supplementary material (App. Tab.11,12) that while significantly lower values in these parameters degrade our performance, the benefit from using larger values saturates.
@@ 7 Limitations
Class-specific performance. In some classes we do not perform as well compared to baseline approaches. A better understanding of the cases where our method fails would be beneficial for deploying it in practice.
@@ 15 Further Image Anomaly Detection Ablation
We include here the ablation tables for the number of random projections and number of bins, for logical anomaly detection (Tab.11,12).
@@ 16 Time Series Anomaly Detection Ablations
Effect of number of pyramid levels and window size. We ablate the two hyperparameters of the time-series feature extraction: the number of pyramid windows used $L$, and the number of samples per window $\tau$ (see Sec.4.2). We find that in both cases the results are not sensitive to the chosen parameters (Tab.15,16).